# Minimum Strict Consistent Subset in Paths, Spiders, Combs and Trees

### v2

@@ 4.4 Running Time Analysis
One might wonder that for any fixed pair $(x,y)$ if the size of solutions of $T(x,y,z)$, i.e. $S(x,y,z)$, is monotonic (increasing or decreasing) with respect to $\mathrm{d}(x,z)$ then we could save an $\mathcal{O}(n)$ factor in the computation of $\mathrm{T}(x,y)$. However, this is not always the case; Figure 6 shows that if we increase $\mathrm{d}(x,z)$ then $S(x,y,z)$ could increase (Figure 6(A)) or decrease (Figure 6(B)).
Figure 6: If we increase $\mathrm{d}(x,z)$ then $S(x,y)$ can increase (A) or decreases (B)
@@ 5.1 Algorithm for Finding MSCS in Paths
We employ arguments that are quite similar to those of the paper [8], which addressed the algorithm of MCS in paths. As a result, the algorithm for MSCS in paths is as follows.
@@ 5.2 Algorithm for Finding MSCS in Spiders
Let $B$ denote the block containing the center of $\mathrm{T}$. Any solution must contain some vertex from $B$, among which one is closest to the center. We examine each vertex $b\in B$ as a vertex in the solution that is closest to the center. Then we break the tree into some paths each of which is defined by two endpoints that are $b$ and a leaf of $\mathrm{T}$. We solve the MSCS problem on each path independently (with the constraint that $b$ must be in the solution) and then take the union of the solutions as a solution for the MSCS problem on $T$. The total running time is $\lvert B\rvert\cdot\mathcal{O}(n)=\mathcal{O}(n^{2})$.
@@ 6 Concluding Remarks
As the paper MSCS is a variant of the paper MCS, therefore, the algorithms for MSCS problems in other graph classes are open. Moreover, as MSCS is NP-hard, so the approximation algorithm, FPT, APX-hard, PTAS are also open for general graphs.

### v3

@@ 5 Algorithm for Finding MSCS in Paths, Spiders, Combs
The definitions of the paths, spiders and combs must be found in [6] and [8]. Therefore, the algorithms in paths, spiders, and combs are as follows:
@@ 5.2 Algorithm for Finding MSCS in Spiders
Let $B$ denote the block containing the center of the spider graph. Any solution must contain at least one vertex from $B$ according to Lemma 3.1, among which one is closest to the center. We examine each vertex $b\in B$ as a vertex in the solution that is closest to the center. Then, we break the spider into some paths, each of which is defined by two endpoints $b$ and the end vertex of the leg of the spider. That is, for each leg of the spider, we take the end vertex of the leg and the vertex $b$, which forms a path. We solve the MSCS problem on each path independently (with the constraint that $b$ must be in the solution) and then take the union of the solutions as a solution for the MSCS problem in the spider. The total running time is $\lvert B\rvert\cdot\mathcal{O}(n)=\mathcal{O}(n^{2})$.
@@ 6 Concluding Remarks
The problem MSCS is a variant of the MCS. Hence, the algorithms for MSCS problems in other graph classes are open, especially for planar graphs. Furthermore, as we demonstrate that MSCS is NP-hard in general graphs, the approximation algorithm, FPT, APX-hard, and PTAS, are also open for general graphs.